# Dynamics of Clusters of Anyons in Fractional Quantum Hall Fluids

### v1

@@ Abstract
We investigates the effective interactions between anyons emerging from either model or realistic bare electron-electron (e-e) interactions in Laughlin and Moore-Read fractional quantum Hall (FQH) fluids. Instead of being purely repulsive or attractive, such anyons display rich dynamics with interesting experimental consequences. Two Laughlin anyons prefer to form bound states with short-range e-e interactions, leading to $32e/$ 3 bunched quasiholes at low temperatures instead of the $e/3$ quasiholes. In non-Abelian Moore-Read FQH phases, two $e/4$ quasiholes can fuse into topologically distinct “1” or “$\psi$” anyons that are no longer degenerate with realistic two-body e-e interactions. This suggests the possibility to energetically separate and manipulate the two types of anyons by tuning the bare electron-electron interactions. We propose that the recently developed high-resolution STM measurements can be used to probe effective anyon interactions when anyons are clustered together after the tunneling of electrons. The local density of states from various bound states of anyon clusters are simulated for both Abelian and non-Abelian systems with (screened) Coulomb interactions.
@@ I Introduction
Anyons with fractional charges and fractional statistics are particles only existing in two-dimensional manifolds because of the nontrivial braid group in 2D Wilczek (1982a, b); Arovas et al. (1985); Read and Moore (1991); Wu (1991); Lundholm (2023). The fractional quantum Hall (FQH) states, realized in a two-dimensional electron gas (2DEG) with strong electronic interaction subjected to a perpendicular strong magnetic field at low temperature, provides an ideal platform to study the exotic properties of anyons Arovas et al. (1984); Halperin (1984); Sharma et al. (2021); Scarola et al. (2000). Quasiholes that can be regarded as anyons in the incompressible FQH fluids are low-lying collective excitations of electrons created by adiabatically inserting fluxes into the FQH ground states with distinct topological orders Laughlin (1983); Wen (1995). The simplest example is the Laughlin $1/m$ anyons with fractional statistics $\nu\pi$ and fractional charge $\nu e$, created by inserting one flux into the Laughlin ground state at filling factor $\nu=1/m$ Kjønsberg and Leinaas (1997). More intriguing cases are the non-Abelian phases in which several anyons with non-Abelian statistics are created by one flux insertion. The great interest in non-Abelian anyons is rooted in the possibility of their application to fault-tolerant topological quantum computing Kitaev (2003); Nayak et al. (2008); Field and Simula (2018).
@@
Much effort has been made to understand the properties of emergent quasi-particles in FQH fluids, such as their fractional statistics, the relationship between them in different FQH phases, the size of anyons and their structures, etc Baraban et al. (2009); Wu et al. (2014); Trung et al. (2023); Trung and Yang (2021); Johri et al. (2014); Ji et al. (2024). One modern perspective is that anyons are “elementary particles” within the respective conformal Hilbert spaces (CHSs) defined by model Hamiltonians, or more generally, by local exclusion conditions (LEC) Read and Moore (1991); Girvin and Prange (1987); Simon et al. (2007); Hansson et al. (2017); Yang (2019); Yang and Balram (2021). These CHSs with emergent conformal symmetry are spanned by the ground state and all quasihole states Yang (2022a); Yuzhu and Bo (2023). If only the model Hamiltonian defining the CHS exists as the electronic interaction term, then the anyons are massless (i.e. it costs no energy to create a single anyon) and free (i.e. the anyons are non-interacting) particles. With more realistic electron interactions, they will gain self-energy and start to interact Wójs (2001). Previous studies manifest the hierarchical structure of CHSs, implying the nontrivial internal structures of anyons as the elementary particles in the corresponding CHSs Yang (2022a); Read (2009). These intricate properties and the nature of CHSs serve as the foundation for revealing the exotic properties of different topological FQH phases and their relationship Yang (2022a); Trung and Yang (2021); Yuzhu and Bo (2023); Wang and Yang (2022); Liang et al. (2024).
@@
Experimental progress, such as shot noise, quantum point contact, and anyon interferometry, provides direct detection for both Abelian and non-Abelian natures of certain anyons Bartolomei et al. (2020); Lee and Sim (2022); Ruelle et al. (2023); Lee et al. (2023); Glidic et al. (2023); Banerjee et al. (2018). While these experimental methods often require a dilute anyon gas for braiding and quantum tunneling, meaning a large separation between anyons, other phases such as Wigner crystal, stripes, and bubbles will compete with FQH phases when the anyon’s density is high Levesque et al. (1984); MacDonald and Murray (1985); Samkharadze et al. (2016); Graß et al. (2018); Wang (2023). This leads to the rich phase diagrams of 2DEG and makes it interesting to understand the dynamics of a cluster of anyons when they are close and their interaction becomes important. Recent advancements in atomic-scale spectroscopy within ultra-clean, diverse 2D systems have enabled the use of local scanning tunneling microscopy (STM) to measure tunneling processes of a cluster of anyons, probe anyon clusters’s dynamics and provide valuable insights into their energies and fusion rules Papić et al. (2018); Feldman and Halperin (2021); Coissard et al. (2022); Hu et al. (2023). In such experiments, the proximity of anyons amplifies the role of effective interactions that strongly influences their behavior Gattu et al. (2024).
@@
In this work, we focus on the effective microscopic interaction between anyons in the Laughlin and Moore-Read (MR) phases, derived from both the model and the realistic bare interactions between electrons. Such effective interactions are highly nontrivial, reflecting information about the particle statisticsArovas et al. (1985, 1984) and the spatial extension of particle sizes (i.e., anyons are not point particles)Yang (2022b); Ji et al. (2024). This leads to the formation of few-anyon clusters and are potentially important for many-anyon quantum fluids with both universal and non-universal dynamics. In particular, we show that two Laughlin 1/3 anyons form a bound “molecule” with short-range electron-electron (e-e) interactions, providing a possible explanation for the anomalous results in quantum point contact experiments of Laughlin phases Bid et al. (2009); Ferraro et al. (2010); Carrega et al. (2011); Sabo et al. (2017). Moreover, different dynamics of non-Abelian $\psi$-type and Abelian $1$-type anyons in the Moore-Read phase with the same e-e interactions break the ground state degeneracy, implying that the Majorana fermion is no longer massless in the presence of the realistic interactions. For both Abelian and non-Abelian anyons, we demonstrate that the local density of states (LDOS) detected by high-resolution STM depends on the interactions within the anyon clusters created by adding or removing a single electron. This makes STM a powerful tool for probing the effective anyon interactions in topological FQH phases.
@@ II.1 Spherical Geometry and Jack Polynomial Formalism
Numerics in this paper are conducted on a sphere enclosing a Dirac magnetic monopole at the center. The monopole generates a uniform magnetic field normal to the surface of the sphere with total flux $2S$ in units of the magnetic flux quanta $\Phi=h/e$, where $2S$ must be an integer Dirac (1931). Electrons on the sphere can be regarded as holding a spin of $l$, $l=|S|+n$ where $n=0,1,2,...$ is the LL index. Single-particle orbitals in one LL are the eigenstates of the angular momentum operators $\hat{L}^{2}$ and its azimuthal part $\hat{L}_{z}$, indexed by $s=-l,-l+1,...,l$. So a single-particle quantum state can be labeled by |$\ket{l,s}$. For a many-body system, the total angular momentum operator and its components are defined as $\hat{L}^{2}=\hat{L}^{2}_{x}+\hat{L}^{2}_{y}+\hat{L}^{2}_{z}$, $\hat{L}_{\alpha}=\sum_{i}\hat{L}_{\alpha,i}$ where $\alpha=x,y,z$ and $i$ is the particle index. The total angular momentum of two electrons can be any integer ranging from 00 to $2l$.
Assuming the electrons are fully spin-polarized, the interaction Hamiltonian in $n_{th}$ LL without LL mixing can be written as
$\hat{H}_{\text{int}}=\frac{1}{2}\sum_{s_{1}+s_{2}\atop=s_{3}+s_{4}}\bra{s_{1},s_{2}}\hat{V}^{(n)}\ket{s_{3},s_{4}}\hat{c}_{s_{1}}^{\dagger}\hat{c}_{s_{2}}^{\dagger}\hat{c}_{s_{3}}\hat{c}_{s_{4}}$ (1)
where $\hat{c}_{s_{r}}^{\dagger}$ ($\hat{c}_{s_{r}}$) is creation (annihilation) operator to the $s_{r}$ orbital with $r$ being the particle index. $\hat{V}^{(n)}$ is the electronic interaction projected to $n_{th}$ LL. $l_{B}=\sqrt{\hbar/eB}$ is the magnetic length, serving as the fundamental length scale in quantum Hall (QH) systems under a perpendicular magnetic field whose magnitude is $B$. Eq.(1) is challenging to handle directly for a generic interaction. But any $\hat{V}^{(n)}$ can be decomposed into a linear combination of Haldane pseudopotentials (PPs) asHaldane (1983); Fano et al. (1986):
@@
$\hat{V}_{m}$ is the pseudopotential that projects a general state to the subspace in which the relative angular momentum between any two particles is $m$. The decomposition coefficient $c_{m}^{(n)}$ represents the energy cost for a pair of electrons to have relative angular momentum $m$. $m$ is constrained to be odd (even) for fermionic (bosonic) FQH systems due to symmetry requirementHsiao (2020); Zhu et al. (2015). The electron pair’s relative angular momentum $m$ and the total angular momentum $L$ satisfies $m+L=2l$. Unless otherwise noted, all the analyses in this paper will be conducted in the lowest LL (LLL), as the case under a strong magnetic field, and the LL index $n$ is dropped hereafter. Larger $m$ corresponds to longer-range interaction, which is evident from the two-particle wavefunctionHaldane (1983); Greiter (2011)
where $u$ and $v$ are spinor coordinates on the sphere. The angular separation of two particles on this sphere is $|u_{1}v_{2}-u_{2}v_{1}|$. A larger $m$ effectively enhances the amplitude of configurations with greater spatial separation.
The advantage of PPs is that it provides an intuitive physical picture and sparsifies the Hamiltonian matrix, thereby simplifying numerical calculations. More importantly, they define model Hamiltonians whose eigenstates are the model wavefunctions of FQH phases. For instance, the eigenstates of the model Hamiltonian $\sum_{m}\hat{V}_{2m-1}$ ($m=1,2,3,...$) are the Laughlin model wavefunctions, describing both the ground states and anyonic quasihole states at the filling factor $\nu=1/(2m+1)$. These states constitute the null space of the model Hamiltonian $\sum_{m}\hat{V}_{2m-1}$ and span the corresponding conformal Hilbert space (CHS). Understanding the effects of individual PPs can also help us predict the influence of realistic interactions that are linear combinations of PPs.
Jack polynomial formalism expresses the many-body wavefunctions of FQH states by root configurations with different admissible rulesBernevig and Haldane (2008a, b). Root configurations consist of “1” or “0”, each representing a single-particle orbital on the sphere with azimuthal angular momentum $s=-l,-l+1,...,l$ from the leftmost (north pole of the sphere) to the rightmost (south pole of the sphere). “1” indicates an orbital occupied by one electron, whereas “0” represents an unoccupied orbital. Laughlin 1/3 states satisfy (1,3) admissibility rule, meaning that there is no more than 1 electron in 3 consecutive orbitals. The ground state for Laughlin 1/3 conformal Hilbert space $\mathcal{H_{L}}$ is the highest density stateYang (2021):
$1001001...001001$ (4)
@@
[…]
@@ II.2 Dynamics of Two Laughlin 1/3 Anyons
with $t\gg 1$, projecting the studied Hilbert space to $\mathcal{H_{L}}$ (null space of $\hat{V}_{1}$) where the elementary particles are the Laughlin $1/3$ anyons, undressed from neutral excitations outside of $\mathcal{H_{L}}$. Without $\hat{V}_{2k+1}$, these anyons are massless and non-interacting as an ideal gas in $\mathcal{H_{L}}$. After adding a relatively small $\hat{V}_{2k+1}$, anyons will gain mass and start to interact. Figure 1 (a) illustrates the effective interaction between two Laughlin $1/3$ anyons as a function of $\Delta M$ under the influence of the model electron-electron interactions $\hat{V}_{3}$ and $\hat{V}_{5}$. The interaction energy is defined as the total energy subtracted by the ground-state energy and the creation energies of all the quasiholes.
@@
$E_{g}$ stands for the ground state energy and $E_{qh}$ is the self-energy of one quasihole (energy cost to create one quasihole). The interacting energies between two anyons are negative at a relatively small separation under both $\hat{V}_{3}$ and $\hat{V}_{5}$ electron-electron interactions in $\mathcal{H_{L}}$. Consequently, the two anyons will combine to form molecule-like structures with $\Delta M=2$ or $4$ being the lowest energy states respectively. This behavior is likely due to the non-trivial statistics and real-space density oscillations of the anyons, which can both introduce unusual effective interactions of anyons Laughlin (1988); Ji et al. (2024); Trung et al. (2023).
Figure 2: Energy spectrum of three-quasihole Laughlin 1/3 state with electron-electron (a) Coulomb interaction (b) Yukawa interaction when $\lambda=1/l_{B}$, both in the LLL; (c) Energy differences between the $\Delta M=6$ state and its competing ones-$\Delta M=8$ (olive circles) and $\Delta M=12$ states (green squares)-plotted as a function of $\lambda$, both exhibiting monotonic behaviours. The system size is $N_{e}=10$.
Building on the study with model Hamiltonians, we further explore the impact of realistic e-e interactions on two-anyon clusters. Two Laughlin $1/3$ anyons tend to separate with Coulomb interaction, shown in Figure 1 (b). Forming an anyon molecule requires a relatively large coefficient for $\hat{V}_{1}$ and the dominance of $\hat{V}_{2k+1}$ with small $k$ based on the results of model Hamiltonians. We hence examine the anyon dynamics under the short-range Yukawa interaction that is commonly used to incorporate the screening effect theoretically.
@@
where $\lambda$ is the inverse of coherent length in units of $1/l_{B}$. As $\lambda$ increases, the lowest energy state of the two-anyon cluster undergoes a monotonic transition from a deconfined state—as in Coulomb interaction’s case—to a bunching state with $\Delta M=2$, e.g. the green dashed line in Figure 1 (b) manifesting the effective anyon interaction induced by $\lambda=0.375$ Yukawa interaction. The inset of Figure 1 (b) shows the monotonic trend of this transition, that the energy differences between the bunching $\Delta M=2$ state and the separated state decrease gradually as $\lambda$ increases, changing from positive to negative.
The two-anyon “molecule” introduced by the short-range electronic interactions provides a plausible explanation for the unexpected results in quantum point contact experiments, where the effective charge carrier shows a temperature dependence, transitioning from $32e/$ 3 at low temperatures ($\sim 10mK$) to $e/3$ at higher temperatures ($\sim 100mK$) Bid et al. (2009); Ferraro et al. (2010); Carrega et al. (2011); Sabo et al. (2017). In shot noise experiments, Coulomb screening will arise from the 2D dielectric environment, nearby metallic gates, and quantum point contact potentials, which will distort electric fields, suppress long-range interactions, and enhance short-range effects Golub (2007); Balram et al. (2013); Spånslätt et al. (2020). Based on the above discussion, at near absolute zero temperature with the effect of short-range interaction, the effective charge carriers in the quantum fluid are bunching “molecules” with $e^{*}=2e/3$ formed by two Laughlin 1/3 anyons. Thermal fluctuations will gradually break apart these “molecules” as temperature rises, and the charge carrier will completely change back to the single Laughlin 1/3 anyons with $e^{*}=e/3$ in the end.

### v2

# Dynamics of Anyon Clusters in Fractional Quantum Hall Fluids

@@ Abstract
In fractional quantum Hall fluids, the quasiparticle excitations are anyons with fractional charges and statistics. Effective interactions among the anyons can be induced by either model or realistic electron-electron (e-e) interactions. Without losing the generality, we investigate such phenomena for the Laughlin $1/3$ and Moore-Read (MR) non-Abelian phases. Anyons display rich internal dynamics in both cases that will lead to interesting experimental consequences. In particular, bound states of two Laughlin anyons are preferred under short-range e-e interactions, leading to $2e/3$ – instead of $e/3$ – effective charge carriers at low temperatures, which have been seen in several experiments. MR phases host two topologically distinct fusion channels: “1” and $\psi$”. The different effective interactions of $e/4$ anyons in the two sectors suggest the vanishing of the degeneracy of fusion channels when the e-e interaction is no longer its model Hamiltonian, in which case different bound states could also appear. This indicates the possibility of energetically manipulating the two types of anyons by tuning the bare e-e interactions. We point out that the results of recently developed high-resolution STM measurements will be affected by the effective anyon interactions, where anyons are clustered together after the tunneling of electrons. The low-lying parts of the local density of states affected by various anyon clusters are simulated for both Abelian and non-Abelian systems with (screened) Coulomb interactions.
@@ I Introduction
Anyons with fractional charges and fractional statistics are particles only existing in two-dimensional manifolds because of the nontrivial braid group in 2D [1, 2, 3, 4, 5, 6]. The fractional quantum Hall (FQH) states, realized in a two-dimensional electron gas (2DEG) with strong electronic interaction subjected to a perpendicular strong magnetic field at low temperature, provide an ideal platform to study the exotic properties of anyons [7, 8, 9, 10]. Quasiholes that can be regarded as anyons in the incompressible FQH fluids are low-lying collective excitations of electrons created by adiabatically inserting fluxes into the FQH ground states with distinct topological orders [11, 12]. The simplest example is the Laughlin $1/m$ anyons with fractional statistics $\nu\pi$ and fractional charge $\nu e$, created by inserting one flux into the Laughlin ground state at filling factor $\nu=1/m$ [13]. More intriguing cases are the non-Abelian phases in which several anyons with non-Abelian statistics are created by one flux insertion. The great interest in non-Abelian anyons is rooted in the possibility of their application to fault-tolerant topological quantum computing [14, 15, 16].
@@
Much effort has been made to understand the properties of emergent quasi-particles in FQH fluids, such as their fractional statistics, the relationship between them in different FQH phases, the size of anyons and their structures, etc [17, 18, 19, 20, 21, 22]. One modern perspective is that anyons are “elementary particles” within the respective conformal Hilbert spaces (CHSs) defined by model Hamiltonians, or more generally, by local exclusion conditions (LEC) [4, 23, 24, 25, 26, 27]. These CHSs with emergent conformal symmetry are spanned by the ground state and all quasihole states [28, 29]. If only the model Hamiltonian defining the CHS exists as the electronic interaction term, then the anyons are massless (i.e., it costs no energy to create a single anyon) and free (i.e., the anyons are non-interacting) particles. With more realistic electron interactions, they will gain self-energy and start to interact [30]. Previous studies manifest the hierarchical structure of CHSs, implying the nontrivial internal structures of anyons as the elementary particles in the corresponding CHSs [28, 31]. These intricate properties and the nature of CHSs serve as the foundation for revealing the exotic properties of different topological FQH phases and their relationship [28, 20, 29, 32, 33].
@@
Experimental progress, such as shot noise, quantum point contact, and anyon interferometry, provides direct detection for both Abelian and non-Abelian natures of certain anyons [34, 35, 36, 37, 38, 39, 40]. While these experimental methods often require a dilute anyon gas for braiding and quantum tunneling, meaning a large separation between anyons, other phases such as Wigner crystal, stripes, and bubbles will compete with FQH phases when the anyon’s density is high [41, 42, 43, 44, 45]. This leads to the rich phase diagrams of 2DEG and makes it interesting to understand the dynamics of a cluster of anyons when they are close and their interaction becomes important. Recent advancements in atomic-scale spectroscopy within ultra-clean, diverse 2D systems have enabled the use of local scanning tunneling microscopy (STM) to measure tunneling processes of a cluster of anyons, probe anyon clusters’ dynamics, and provide valuable insights into their energies and fusion rules [46, 47, 48, 49]. In such experiments, the proximity of anyons amplifies the role of effective interactions that strongly influence their behavior [50].
@@
In this work, we focus on the effective microscopic interaction between anyons in the Laughlin and Moore-Read (MR) phases, derived from both the model and the realistic bare interactions between electrons. Such effective interactions are highly nontrivial, reflecting information about the particle statistics[3, 7] and the spatial extension of particle sizes (i.e., anyons are not point particles)[51, 22]. This leads to the formation of few-anyon clusters and are potentially important for many-anyon quantum fluids with both universal and non-universal dynamics. In particular, we show that two Laughlin 1/3 anyons form a bound “molecule” with short-range electron-electron (e-e) interactions, providing a possible explanation for the anomalous results in quantum point contact experiments of Laughlin phases [52, 53, 54, 55]. Moreover, different dynamics of non-Abelian $\psi$-type and Abelian $1$-type anyons in the Moore-Read phase with the same e-e interactions break the ground state degeneracy, implying that the Majorana fermion is no longer massless in the presence of the realistic interactions. Here the term ”dynamics” is used in the general sense for the parametric evolution of quasihole clusters’ energies, which will manifest themselves in various experiments where dynamical properties of the quantum Hall states are measured. For both Abelian and non-Abelian anyons, we demonstrate that the local density of states (LDOS) detected by high-resolution STM depends on the interactions within the anyon clusters created by adding or removing a single electron. This makes STM a powerful tool for probing the effective anyon interactions in topological FQH phases.
@@ II.1 Spherical Geometry and Jack Polynomial Formalism
[…]
Assuming the electrons are fully spin-polarized, an isotropic two-body interaction in the $n_{th}$ LL, neglecting LL mixing, can be written as
[…]
The projected interaction can then be decomposed as:
@@
[…]
$|\hat{V}_{m}=\sum_{M=-L}^{L}\ket{L,M}\bra{L,M}$ (4)
is the PP operator, which projects any two-particle state onto the subspace of relative angular momentum $m$. Unless otherwise stated, all analyses in this paper are performed in the lowest LL (LLL), corresponding to the strong magnetic field limit, and the LL index $n$ is omitted hereafter. PPs provide both an intuitive physical picture and a sparse representation of the interaction, thereby simplifying numerical calculations. More importantly, they define model Hamiltonians whose eigenstates are the model wavefunctions of certain FQH phases. For instance, the eigenstates of the model Hamiltonian $\sum_{k}^{p}\hat{V}_{2k-1}$, $k=1,2,$..., are the fermionic Laughlin model wavefunctions, describing both the ground states and anyonic quasihole states at the filling factor $\nu=1/(2p+1)$. These states constitute the null space of the model Hamiltonian $\sum_{k}^{p}\hat{V}_{2k-1}$ and span the corresponding conformal Hilbert space (CHS). Understanding the effects of individual PPs can also help us predict the influence of realistic interactions that are linear combinations of PPs.
Jack polynomial formalism expresses the many-body wavefunctions of FQH states by root configurations with different admissible rules[62, 63]. Root configurations consist of “1” or “0”, each representing a single-particle orbital on the sphere with azimuthal angular momentum s=l,l+1,,ls=-l,-l+1,...,l from the leftmost (north pole of the sphere) to the rightmost (south pole of the sphere). “1” indicates an orbital occupied by one electron, whereas “0” represents an unoccupied orbital. Laughlin 1/3 states satisfy (1,3) admissibility rule, meaning that there is no more than 1 electron in 3 consecutive orbitals. The ground state for Laughlin 1/3 conformal Hilbert space $\mathcal{H_{L}}$ is the highest density state[64]:
10010010010011001001...001001 (5)
@@
 0  0 $1\displaystyle\hbox{\hskip 4.16663pt\hskip-2.5pt\hbox{$0$}\hskip-2.5pt\hskip-4.16663pt\raisebox{-5.55556pt}{\hbox{$\mkern 2.0mu\circ$}}\hskip-4.16663pt\hskip 4.16663pt}\hbox{\hskip 4.16663pt\hskip-2.5pt\hbox{$0$}\hskip-2.5pt\hskip-4.16663pt\raisebox{-5.55556pt}{\hbox{$\mkern 2.0mu\circ$}}\hskip-4.16663pt\hskip 4.16663pt}100100$...1001 (6)
@@ II.2 Dynamics of Two Laughlin 1/3 Anyons
with $t\gg 1$, projecting the studied Hilbert space to $\mathcal{H_{L}}$ (null space of $\hat{V}_{1}$) where the elementary particles are the Laughlin $1/3$ anyons, undressed from neutral excitations outside of $\mathcal{H_{L}}$. In our calculations, $t=100$ is typically set, which is sufficient to numerically enforce the projection, shown by a finite energy gap between the quasihole states and all the other states above them in the spectrum. In fact, we find that a value of $t$ just one order of magnitude larger than the coefficient of the additional e-e interaction term already ensures an effective separation between the quasihole sector and the higher-energy continuum.
Without $\hat{V}_{2k+1}$, these anyons are massless and non-interacting as an ideal gas in $\mathcal{H_{L}}$. After adding a relatively small $\hat{V}_{2k+1}$, anyons will gain mass and start to interact. Figure 1 (a) illustrates the effective interaction between two Laughlin $1/3$ anyons as a function of $\Delta M$ under the influence of the model electron-electron interactions $\hat{V}_{3}$ and $\hat{V}_{5}$. The interaction energy is defined as the total energy subtracted by the ground-state energy and the creation energies of all the quasiholes.
@@
$E_{g}$ stands for the ground state energy and $E_{qh}$ is the self-energy of one quasihole (energy cost to create one quasihole). To compute $E_{qh}$, we introduce one additional flux into the system while keeping the number of electrons fixed, and subtract the resulting ground-state energy from $E_{g}$. This defines the creation energy of a single, well-isolated quasihole, independent of any angular momentum, spatial configuration, or presence of other quasiholes.
The interacting energies between two anyons are negative at a relatively small separation under both $\hat{V}_{3}$ and $\hat{V}_{5}$ electron-electron interactions in $\mathcal{H_{L}}$, shown in Figure 1. Consequently, the two anyons will combine to form molecule-like structures with $\Delta M=2$ or $4$ being the lowest energy states, respectively. Finite-size scalings in (c, d) demonstrate that the key features, such as the formation and stability of the bound anyon cluster state, are robust and converge well when increasing system size. While a detailed understanding of the effective interaction’s origin remains to be established, our studies and several indirect arguments in the literature support the relevance to the non-trivial statistics and real-space density oscillations of the anyons [65, 22, 19]. A deeper investigation into the precise roles of these features is an open question for future work.
Figure 2: Energy spectrum of three-quasihole Laughlin 1/3 state with electron-electron (a) Coulomb interaction (b) Yukawa interaction when $\lambda=1/l_{B}$, both in the LLL; (c) Energy differences between the $\Delta M=6$ state and its competing ones-$\Delta M=8$ (olive circles) and $\Delta M=12$ states (green squares)-plotted as a function of $\lambda$, both exhibiting monotonic behaviours. Red crosses indicate quasihole states, blue pluses denote other excitations outside the QH manifold. The system size is $N_{e}=10$.
Building on the study with model Hamiltonians, we further explore the impact of realistic e-e interactions on two-anyon clusters. Forming an anyon molecule requires a relatively large coefficient for $\hat{V}_{1}$ and the dominance of $\hat{V}_{2k+1}$ with small $k$ based on the results of model Hamiltonians. We hence examine the anyon dynamics under the short-range Yukawa interaction that is commonly used to incorporate the screening effect theoretically:
@@
where $\lambda$ is the inverse of coherent length in units of $1/l_{B}$. Two Laughlin $1/3$ anyons tend to separate with Coulomb interaction, shown in Figure 1 (b). As $\lambda$ increases, the lowest energy state of the two-anyon cluster undergoes a monotonic transition from a deconfined state to a bunching state with $\Delta M=2$, such as the green dashed line manifested where the effective anyon interaction is induced by $\lambda=0.375$ Yukawa interaction. The inset of Figure 1 (b) shows the monotonic trend of this transition, where the energy differences between the bunching $\Delta M=2$ state and the separated state decrease gradually as $\lambda$ increases, changing from positive to negative. Since screening uniformly reduces the overall energy scale, we normalize the pseudopotential decomposition by fixing the coefficient of $\hat{V}_{3}$ to unity, and assign the remaining pseudopotential coefficients according to their original relative ratios. This ensures that all quantities such as energy gaps and total energies remain comparable across different values of $\lambda$ (and Coulomb).
The two-anyon “molecule” introduced by the short-range electronic interactions provides a plausible explanation for the unexpected results in quantum point contact experiments, where the effective charge carrier shows a temperature dependence, transitioning from $2e/3$ at low temperatures ($\sim 10mK$) to $e/3$ at higher temperatures ($\sim 100mK$) [52, 53, 54, 55]. In shot noise experiments, Coulomb screening will arise from the 2D dielectric environment, nearby metallic gates, and quantum point contact potentials, which will distort electric fields, suppress long-range interactions, and enhance short-range effects [66, 67, 68]. Based on the above discussion, at near absolute zero temperature with the effect of short-range interaction, the effective charge carriers in the quantum fluid are bunching “molecules” with $e^{*}=2e/3$ formed by two Laughlin 1/3 anyons. Thermal fluctuations will gradually break apart these “molecules” as temperature rises, and the charge carrier will completely change back to the single Laughlin 1/3 anyons with $e^{*}=e/3$ in the end. A quantitative comparison between the binding energy of the anyon “molecule” under Yukawa interaction with $\lambda=0.375/l_{B}$ ($10^{-24}\sim 10^{-25}\text{J}$) and the thermal energy at $T\sim 10-100\text{mK}$ confirms that their scales are comparable, supporting this scenario.